OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 0.07k+0.15(200)=0.09(K+200) I Need to show all work. I Am Horrible At Math Can Someone Please Take Me Through The Steps to get 600 For The Answer

jimthompson5910 (jim_thompson5910):

Start by multiplying both sides by 100 0.07k+0.15(200)=0.09(K+200) 100(0.07k+0.15(200))=100*0.09(K+200) 7k + 15(200) = 9(k+200) what's next?

OpenStudy (anonymous):

distribute the 9(k+200) Right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that gives you what

OpenStudy (anonymous):

7k+15(200)=9k+200

jimthompson5910 (jim_thompson5910):

you have to multiply that 9 by EVERY term inside the parenthesis

jimthompson5910 (jim_thompson5910):

so instead of 200, it would be 9*200 = 1800

OpenStudy (anonymous):

k=-600